Solve the following equation by completing the square. (Enter solutions from smallest to largest. If there is one solution with a multiplicity of two, enter it in both answer boxes.)
x^2+4x=12
x = ?
x = ?
Please show me the steps to get the answer. Thanks

Answer by jorel1380(3719)

x2+4x=12
x2+4x-12=0
(x+6)(x-2)=0
x=-6 or 2..